Greetings everyone


I need of your help. Converting this basic code into matlab code in order to complete the table below please. Thank you. I will give a good rating
100 J1 110 N = N + 1 120 IF RND > N/365 THEN 110 130 MM+N+1 140 PRINT J, N +1, M/3 150 N- 0 160 JJ 1 170 GOTO 110 Before running the simulation, estimate the long-run average here: Number Questioned in this Experiment Average Number Questioned per Experiment Experiment 4 10 25 50 100 1000 10000
Answer #1

Code in Matlab

N = 0
J = 1
M=0
while(1)
N = N+1
while rand()>N/365
N =N+1
end
M = M + N + 1
disp(J)
disp(N + 1)
disp(M/J)
N =0
J = J + 1
end

According to loop this will be indefinite loop